# Probing the Vulnerability of Large Language Models to Polysemantic Interventions

### v1

@@ 3.2 Steering with Gradient Vector for Token Intervention
In this experiment, to compare the intervention performance between SAE feature directions and gradient vectors, we use the same target and interference features in the previous experiment for tests. Also, the steering vector is scaled within the same range. As shown in Figure 4, using token gradients as steering vectors on GPT-2-Small yielded better results than using feature directions. Intriguingly, when applying token gradient-based steering vectors to Pythia-70M, the relationship between interference scale and intervention effectiveness is flattened. This may be attributed to the fact that token gradients are independent of SAE-defined interference levels, and that SAE features can exhibit a degree of arbitrariness (Paulo & Belrose, 2025; Heap et al., 2025).
Figure 4: Comparing effects of activation engineering via SAE direction and gradients vector. The x-axis follows the same definition as in Figure 3, but attack vectors are derived from token gradients instead of SAE decoder weights. Error bars indicate standard errors. ***, **, and * denote t-test significance at $p<0.001$, $p<0.01$, and $p<0.05$, respectively, compared to the random baseline (i.e., random token gradient directions).
Finding 2: Steering highly activating tokens from distinct but interfering features based on gradient directions can significantly amplify the model’s output on a target feature.
Type Model Intervention Target feature Result
Llama-3.1-8b-Instructed Steering gradient vector: references to the world and its various aspects References to ‘Switzerland’ “I would like to recommend you to spend holidays in” $\uparrow$ Entered $\downarrow$ Dropped Switzerland +0.16 Italy -0.034 Germany +0.089 Greece -0.016 Canada +0.015 Bulgaria -0.012
Llama-3.1-8b-Instructed Prepending the injection text “(placement from placement)” References to locations “In the next weekend we will go to” $\uparrow$ Entered $\downarrow$ Dropped Paris +0.011 another -0.013 * +0.010 H -0.003 - +0.006 K -0.003
@@ 3.3 Prompt Injection for Inference Time Intervention
The number of ways to inject $n$ tokens into a prompt grows super-exponentially with $n$, making optimization costly. Here, we apply a straightforward method by prepending 10 selected tokens to the prompt, allowing their influence to propagate during inference, affect the activation of the target feature, and ultimately impact the model’s output. To assess the generality of our findings, we test two types of target tokens: location names and concepts with strong semantic polarity (e.g., “hate” or “love”) [^4] The rationale for selecting these two features is detailed in Appendix H.1. For each target set, we identify the corresponding SAE features and extract the top-activating tokens from features with either high or low interference. In addition to these, we include two baseline sets for comparison: a random token set and the original target token set. For each prompt, we sample injection tokens 100 times and compute the success rate as the proportion of runs where target-type tokens are elevated into the top-10 predictions. Table 1 presents representative examples and Table 2 shows macro statistics. As shown, high-interference tokens are more effective at elevating target-related tokens into the top-10 predictions than low-interference or random tokens, though still much less effective than directly prepending the target’s tokens.
@@
Locations Pythia-70M $49.74\%$*** $31.05\%$*** $21.76\%$ $22.87\%$
GPT-2-Small $39.92\%$*** $19.05\%$*** $14.74\%$ $13.54\%$
LLaMA3.1-8B-Instruct $47.62\%$*** $27.64\%$*** $20.05\%$ $20.33\%$
Gemma-2-9B-Instruct $39.21\%$*** $26.68\%$*** $15.39\%$ $15.45\%$
Hate/Love Pythia-70M $23.67\%$*** $5.48\%$*** $4.00\%$* $3.39\%$
GPT-2-Small $38.08\%$*** $10.27\%$*** $7.34\%$** $6.23\%$
LLaMA3.1-8B-Instruct $56.08\%$*** $11.24\%$*** $6.44\%$* $5.70\%$
Gemma-2-9B-Instruct $50.79\%$*** $12.27\%$*** $7.34\%$ $7.02\%$
@@
Note: Cell values show the success rate of elevating target-type tokens into the top-10 predictions. Gray-shaded rows indicate black-box interventions. Testing uses a shared token set from the two small models. ***, **, and * denote z-test significance at $p<0.001$, $p<0.01$, and $p<0.05$, respectively, vs. random baseline. High- and low-interference tokens lie in $[0.5,1.0]$ and $[0.2,0.5]$. Details in Appendix H.3.
Finding 3: Injecting top-activating tokens from distinct but interfered features into prompts can significantly amplify the model’s output on a target feature.
3.4 Generalization of Vulnerability
@@
Since token gradient-based and prompt-based interventions do not rely on pre-trained SAEs, they can be applied to models without access to detailed internal representations. To demonstrate this, we target the LLaMA3.1-8B-Instruct and Gemma-2-9B-Instruct models, reusing the two targets from the prompt injection experiment. We focus on the top-activating tokens of features that interfere with the targets, and identify those shared across both Pythia-70M and GPT-2-Small. For example, Pythia-70M contains 1,309 tokens with interference values above $0.5$ for location-related features, while GPT-2-Small has 932 such tokens; their intersection includes 193 shared tokens.
@@
We apply token gradient-based intervention to LLaMA3.1-8B-Instruct and, notably, by extracting steering vectors through random sampling of only one high-interference features’ top five activation texts, we could boost the presence of relevant tokens in the top-10 prediction list with over 95% success rate. Prompt injection interventions are tested on both LLaMA3.1-8B-Instruct and Gemma-2-9B-Instruct. As shown in Table 2, high-interference tokens derived from the two small models can steer both larger models more effectively than random baselines. In hindsight, these results suggest that shared polysemantic structures observed in small models also extend to larger models, indicating generalized vulnerabilities that persist across architectures and training regimes.
Finding 4: The success of black-box attacks from small to large models reveals transferable polysemantic structures and highlights weaknesses in model robustness.
@@ 3.5 Manipulating Activations for Neuron Intervention
To complete our discussion, we also explore models’ vulnerability to interventions on individual neurons. Specifically, we investigate how the degree of polysemanticity in neurons affects the output. For the aggregated features obtained in Section 2.2, we analyze their connections with neurons. Here, we only involve neuron-feature pairs with a connection strength greater than $0.2$. Among the filtered neurons, those connected to only one or two aggregated features account for more than $33\%$ in strongly connected neurons, as shown in Figure 15. In addition to neurons connected to multiple or dozens of aggregated features, there are also some “super-neurons” with connections exceeding $500$. We examine the impact of manipulating these neurons on the model’s output. The experimental results indicate that neurons with higher degrees of polysemanticity are more vulnerable, which means they tend to affect the model’s output more effectively. However, for certain “super-neurons,” the impact on the model is notably asymmetric: masking them results in even less influence than neurons with lower polysemanticity, while amplifying their activations often leads to exponentially greater effects on model behavior.
Figure 5: Effects of activating and deactivating neurons by polysemanticity level. The x-axis shows neuron groups from monosemantic (left) to “super-neurons” (right). “Drop” and “increase” refer to masking and amplifying activations. The y-axis shows the change in weighted cosine similarity. Error bars indicate standard errors.
Finding 5: Neurons with higher polysemanticity have greater influence on model outputs, with “super-neurons” showing an asymmetric effect.
@@ 4 Discussion
In this paper, we investigate the sensitivity of LLMs to structured interventions grounded in their polysemantic representations. Specifically, we examine three types of interventions: (1) feature direction-based, (2) token gradient-based, and (3) prompt-based. Feature direction interventions rely on SAEs. While less effective than gradient-based approaches, they form the basis for deriving token gradient vectors. Token gradient-based interventions are more effective and can be constructed directly from activation texts, without requiring SAE pre-training—though they do assume access to internal activations. Prompt-based interventions require minimal access and, despite their surface-level nature, still yield meaningful behavioral shifts. Additionally, we explore neuron-level interventions, motivated by the uneven distribution of features across neurons. We find that the behavioral impact of masking and amplification correlates with neuron polysemanticity. Notably, we identify a class of “super-neurons,” those encoding over $500$ features, for which amplification significantly alters model behavior, while deactivation results in a markedly reduced effect.
Our findings further show that shared polysemantic structures identified in the two small models can be transferred to guide token- and prompt-level interventions in larger, instruction-tuned black-box models such as LLaMA-3.1-8B-Instruct and Gemma-2-9B-Instruct. This suggests that certain polysemantic features are preserved across architectures and training regimes, exposing a shared representational basis that allows behavioral modulation even without access to internal weights. Notably, this challenges prevailing theories that treat polysemanticity as an incidental artifact of training (Marshall & Kirchner, 2024; Lecomte et al., 2023), which cannot account for the generalization we observe across models. These results raise deeper questions about the nature of polysemanticity in LLMs: are these structures unintended byproducts, or do some of them reflect stable, higher-order organizational patterns? Our findings also strengthen recent evidence of representational consistency and topological stability across models (Huh et al., 2024; Wolfram & Schein, 2025; Lee et al., 2025), even as the origins and functional implications of this consistency remain open challenges.
Our work is among the first to systematically evaluate polysemantic vulnerabilities in LLMs, but it has several limitations, including the range of model sizes tested, intervention depth, and transferability robustness. We discuss these limitations and ethical considerations in Appendix J.
@@ 5 Conclusion
This study systematically investigates the sensitivity of LLMs to structured interventions grounded in the polysemantic representations of two small LLMs. Leveraging SAEs, we show that model behavior can be steered toward specific feature directions by manipulating unrelated but interfering features using three intervention methods. We further show that token- and prompt-based interventions derived from shared polysemantic structures in small models can transfer effectively to larger, black-box models. This points to a stable and transferable polysemantic topology that persists across model architectures and training regimes. Finally, by analyzing the uneven distribution of features across neurons, we examine model sensitivity to neuron-level manipulations across varying degrees of polysemanticity, revealing an asymmetric effect in “super-neurons.” Our findings offer a foundation for future work on the structural properties and representational robustness of LLMs.
@@ Appendix A Related Work
A.1 A Brief Review on LLM Adversarial Attacks
@@ A.2 A Brief Review on SAE-based Intervention Techniques in LLMs
SAE-based interventions represent a promising direction for developing more interpretable and controllable LLMs. Recent research have introduced a diverse set of SAE-based techniques, such as clamping, patching, and causal tracing, applied across a range of use cases (farrell2024applying,cunningham2023sparse,marks2024sparse). Empirical results indicate that these methods can be highly effective. For example, targeted unlearning via SAE features has been shown to suppress undesired capabilities with fewer side effects than global fine-tuning (Khoriaty et al., 2025), while feature-level steering enables more nuanced output control than prompt-based methods alone (Rajamanoharan et al., 2024). A key advantage of SAE-based approaches is their efficiency at inference time: they often require only a forward pass with lightweight vector operations and typically do not require model retraining, making them well-suited for real-time interventions.
@@ Appendix B Impact Statement
This work systematically investigates a semantic vulnerability in LLMs rooted in polysemanticity—where single neurons encode multiple unrelated features. We introduce four complementary approaches that expose this vulnerability: manipulating SAE-derived features, token gradients, and prompts to steer model outputs via semantically unrelated inputs, and intervening at the neuron level to reveal a correlation between polysemanticity and output sensitivity. We also identify a class of “super-neurons” whose amplification disproportionately alters model behavior, while masking them has limited effect. These findings not only highlight the unique characteristics of structural fragility of LLMs but also provide practical tools for probing and controlling their internal mechanisms. Our work lays a foundation for future research in AI safety, enabling both defenses against such vulnerabilities and the development of more efficient, targeted interventions for model alignment and interpretability.
Appendix C Sparse Autoencoder Training
SAEs are a rapidly developing tool for probing the polysemantic structure of neurons (Shu et al., 2025). Given the activation vector $\mathbf{a}\in\mathbb{R}^{d_{\text{embed}}}$ from a particular model layer, an SAE projects it into a higher–dimensional sparse code $\mathbf{f}\in\mathbb{R}^{d_{\text{sae}}}$ in order to disentangle the multiple semantics that a single neuron may simultaneously encode. The forward computation and the resulting feature definition $\mathbf{f}$ are shown below:

### v2

@@ 3.2 Steering with Gradient Vector for Token Intervention
In this section, we treat the top-activating tokens of semantically unrelated yet interfering SAE features as intervention signals (Ferrando et al., 2024). For each interfering feature, we pass its top-activating text through the model and, at the feature’s corresponding layer, compute the gradient of that token with respect to all neurons in that layer. This gradient serves as the steering vector. As shown in Figure 4, using token-gradient steering on both models yields roughly $\sim 10\times$ larger effects than steering along SAE feature directions. Interestingly, token-gradient steering also flattens the relationship between interference scale and intervention effectiveness, and steering the original feature’s gradient direction is less effective than steering the interfering ones. This may stem from the fact that token gradients are not tied to SAE-defined interference levels and SAE features can exhibit a degree of arbitrariness (Paulo & Belrose, 2025; Heap et al., 2025).
Figure 4: Effects of interventions along the gradient direction of interference SAE feature’s top-activating tokens. Subpanels follow the same conventions as Figure 3, but intervention vectors are computed from token gradients rather than SAE decoder weights. For (A–B), error bars indicate 95% confidence intervals; for (C), error bars denote 90%, 95%, and 99% confidence intervals. Results with the alternative metric are shown in Figure 10.
@@ 3.3 Prompt Injection for Inference Time Intervention
In addition to the two intervention methods described above, which directly modify the model’s internal activations, we also investigate whether models are vulnerable to polysemantic interference during inference time. In this case, hidden state manipulation is achieved indirectly by modifying the input prompt. The core idea is that, for a given target feature, we identify highly interfering features, select the tokens with the highest activation values for those features, and inject these tokens into the prompt. This is intended to activate the corresponding features, allowing their influence to propagate and affect the activation of the target feature. The number of ways to inject $n$ tokens into a prompt grows super-exponentially with $n$, making optimization costly. Here, we apply a straightforward method by prepending randomly selected tokens from each interference value range to the prompt. To assess the generality of our findings, we intervene on the model’s production of tokens associated with selected target features, illustrating with three cases: location names, digital numbers, and science-related terms  [^5] More intervention cases are listed in Table 5.. For each target set, we identify the corresponding SAE features and extract the top-activating tokens from features with either high or low interference. In addition to these, we include two baseline sets for comparison: a random token set and the original target token set. For each prompt, we sample injection tokens $100$ times and compute the success rate as the proportion of runs where target-type tokens are elevated into the top-30 predictions. Table 3 presents representative examples and Table 1 shows macro statistics. As shown, in most cases, high-interference tokens are more effective at elevating target-related tokens into the top 30 predictions than low-interference or random tokens, although they are still much less effective than directly prepending the target’s tokens.
@@
Locations Pythia-70M $65.08\%***$ $36.93\%**$ $32.53\%$ $35.06\%$
GPT-2-Small $44.68\%***$ $18.42\%***$ $19.08\%***$ $16.42\%$
Llama-3.1-8B-Instruct $33.84\%***$ $20.78\%***$ $19.63\%*$ $18.24\%$
Gemma-2-9B-Instruct $10.16\%$ $2.71\%*** 11$$1.36\%$ $10.66\%$
Llama-3.1-70B-Instruct $37.23\%***$ $28.21\%***$ $23.09\%**$ $24.48\%$
Number Pythia-70M $65.32\%***$ $35.15\%$ $36.71\%***$ $34.30\%$
GPT-2-Small $55.87\%***$ $30.33\%$ $31.23\%$ $34.71\%$
Llama-3.1-8B-Instruct $55.97\%***$ $31.87\%*$ $32.90\%***$ $30.57\%$
Gemma-2-9B-Instruct $48.42\%***$ $29.93\%***$ $30.64\%***$ $27.16\%$
Llama-3.1-70B-Instruct $25.57\%***$ $8.85\%***$ $6.67\%$ $7.09\%$
Science Pythia-70M $61.66\%***$ $3.13\% 22$$2.78\%$ $28.58\%$
GPT-2-Small $75.70\%***$ $25.93\%***$ $26.07\%***$ $21.71\%$
Llama-3.1-8B-Instruct $49.67\%***$ $20.08\%***$ $18.95\%$ $17.94\%$
Gemma-2-9B-Instruct $46.84\%***$ $20.40\%$ $19.25\%$ $20.15\%$
Llama-3.1-70B-Instruct $67.26\%***$ $48.22\%***$ $43.57\%$ $42.24\%$
@@
Note: Cell values show the success rate of elevating target-type tokens into the top-30 predictions. Gray-shaded rows indicate black-box interventions. Testing uses a shared token set from the two small models. ***, **, and * denote t-test significance at $p<0.001$, $p<0.01$, and $p<0.05$, respectively, vs. random baseline. High- and low-interference tokens lie in $[0.5,1.0]$ and $[0.2,0.5]$ for Pythia-70M, while $[0.3,1.0]$ and $[0.2,0.3]$ in GPT-2-Small. Details in Appendix K.2.
3.4 Generalization of Polysemantic Intervention Vulnerability
@@
Because token-gradient and prompt-based interventions do not rely on SAEs, they can be applied to models without internal access. We therefore target Llama-3.1-8B/70B-Instruct and Gemma-2-9B-Instruct, selecting three target types from the prompt-injection experiment. We still focus on the top-activating tokens of features that interfere with those targets, while retaining tokens extracted from Pythia-70M and GPT-2-Small. Details are illustrated in Appendix K.2
@@
On Llama-3.1-8B-Instruct and Gemma-2-9B-Instruct, we apply gradient-based intervention by extracting steering vectors through random sampling of only one high-interference features’ top five activation texts, and could notably boost the presence of relevant tokens in the top-10 prediction list with over $95\%$ success rate. Prompt injection interventions are tested on Llama-3.1-8/70B-Instruct and Gemma-2-9B-Instruct. As shown in Table 1, high-interference tokens derived from the two small models can steer larger models more effectively than random baselines. In hindsight, these results suggest that shared polysemantic structures observed in small models also extend to larger models, indicating generalized vulnerabilities that persist across architectures and training regimes.
3.5 Analyzing Shared but Counterintuitive Polysemantic Structure
We observe that many similar feature pairs in GPT-2-Small and Pythia-70M that lie far apart in the first-order symbolic manifold, $\mathcal{M}$ (by surface/gloss semantics), nevertheless lie close in both models’ activation space $\mathcal{F}$. To understand this transferability, we consider mechanisms that couple features via higher-order semantic structure, including semantic-priming-type associations (e.g., thematic/scripts, causal, frame roles) and morphological relatedness, even when overt meanings appear unrelated (Mandera et al., 2017; Bojanowski et al., 2017).
To probe these links, we conduct large-scale annotation with DeepSeek-V3 and GPT-5-mini. Models are prompted to identify higher-order semantic relations for shared interfering feature pairs that satisfy three criteria: interference >0.4>0.4, semantic similarity in $\mathcal{M}<0.2$, and cross-model feature-pair semantic similarity >0.5>0.5. In total, 459,229459{,}229 pairs are annotated (prompts, rubric, and head-to-head annotator comparison in Appendix M). Only $27.7\%$ of pairs are judged by at least one model to exhibit a plausible higher-order association, and most such links remain counterintuitive upon post-hoc inspection. Because the association-check task can be challenging for LLMs, we also run a conservative paired-choice evaluation. We sample 3,8003{,}800 high-interference, low-similarity pairs. For each, we randomly draw a comparison pair strictly matched on semantic similarity but with substantially lower interference, then ask GPT-5-mini to select which pair is more related (details in Appendix M). The high-interference pair is chosen $64.3\%$ of the time (Wilson $95\%$ CI [0.628, 0.658][0.628,\,0.658]); a one-sided exact binomial test against $0.5$ is significant ($p<0.001$), corresponding to a log-odds of $0.589$ in favor of the high-interference pair, indicating a statistically meaningful yet modest effect. Taken together, these results both validate our filtering strategy for isolating unrelated pairs in the intervention analyses and point to a striking regularity: LLMs instantiate stable, cross-model polysemantic organization that is often opaque to semantic intuition.
In Appendix M, Table LABEL:tab:semrels reports examples where models detect latent associations between feature pairs and where they do not. One notable case is the last, asterisked example in Table LABEL:tab:semrels: both annotators label it as negative. In our follow-up analysis, however, we hypothesize a biographical–affective link: mentions of “Beethoven” may co-occur with expressions of frustration/suffering, given his late-life deafness and celebrated late-period compositions. These examples suggest that LLM polysemanticity may approximate latent knowledge structures and offer testable hypotheses. In summary, it is possible that human studies could reveal comparable human recognition of the same “weak signals” picked up across models, even if they cannot recall or justify them. A full adjudication of this possibility lies beyond our present scope and we leave for future work.
@@ 3.6 Manipulating Activations for Neuron Intervention
To complete our discussion, we explore models’ vulnerability to interventions on individual neurons. Specifically, we investigate how the degree of polysemanticity in neurons affects the output. For aggregated features obtained through agglomerative clustering under the cosine similarity threshold of $0.4$, we quantify each neuron’s connected number of features. Here, we only involve neuron-feature pairs with a connection strength greater than $0.2$. Among all neurons, those connected to only one or two aggregated features account for more than $33\%$ in strongly connected neurons, as shown in Figure 15. In addition to neurons connected to multiple or dozens of aggregate features, there are also some “super-neurons” with connections exceeding $500$. We examine the impact of manipulating these neurons on the model’s output. Experimental results indicate that neurons with higher degrees of polysemanticity are more vulnerable, which means they tend to affect model outputs more effectively. For certain “super-neurons,” however, the impact on the model is notably asymmetric: masking them results in less influence than neurons with lower polysemanticity, while amplifying their activations often leads to exponentially greater effects on model behavior.
Figure 5: Effects of activating and deactivating neurons by polysemantic level. The $x$-axis indicates neuron categories grouped by the number of connected features (after clustering). The $y$-axis reports the change in weighted cosine similarity. Each box plot centers on the median ($50\%$) and progressively splits the remaining data in half at each level. A and B correspond to GPT-2-Small; C and D correspond to Pythia-70M. A and C show the effect of masking neuron activations, while B and D show the effect of amplifying them.
@@ 4 Discussion
This work makes two contributions. First, we systematically investigate the vulnerability of LLMs to structured interventions grounded in their polysemantic representations. Specifically, we examine three types of intervention: (1) feature direction-based, (2) token gradient-based, and (3) prompt-based. Feature direction interventions rely on SAEs. While less effective than gradient-based approaches, they form the basis for deriving token gradient vectors. Token gradient-based interventions are most effective and can be constructed directly from activation texts, without requiring SAE pre-training—although they assume access to internal activations. Prompt-based interventions require minimal access and, despite their surface-level nature, still yield meaningful behavioral shifts. Additionally, we explore neuron-level interventions, motivated by the uneven distribution of features across neurons. We find that the behavioral impact of masking and amplification correlates with neuron polysemanticity. We also identify a class of “super-neurons,” those encoding over $500$ features, for which amplification significantly alters model behavior, while deactivation results in a markedly reduced effect.
The second contribution lies in our finding that polysemantic, structural vulnerabilities identified in two small models transfer to larger, instruction-tuned black-box models (e.g., Llama-3.1-8B/70B-Instruct, Gemma-2-9B-Instruct) via token-gradient- and prompt-level manipulations, producing predictable behavioral shifts. This suggests that certain polysemantic structures are preserved across architectures and training regimes, exposing a shared representational basis. This directly challenges prevailing theories that treat polysemanticity as an incidental artifact of training (Marshall & Kirchner, 2024; Lecomte et al., 2023). Our exploratory analyses further suggest that these transferable polysemantic structures are not reducible to higher-order relations readily intelligible to humans; we therefore treat these counterintuitive regularities as testable hypotheses about latent knowledge structure. These results sharpen a central question about LLM polysemanticity: are they unintended byproducts or stable, higher-order patterns that await rigorous examination? Our results point towards the latter. Finally, our findings strengthen recent evidence of representational consistency and topological stability across models (Huh et al., 2024; Wolfram & Schein, 2025; Lee et al., 2025), even as the origins and functional implications of this consistency remain open challenges. Our work is the first to systematically evaluate polysemantic structural vulnerabilities in real-world LLMs, but it has several limitations, including intervention depth and transferability robustness tests. We discuss these limitations and ethical considerations in Appendix O.
@@ 5 Conclusion
We systematically probe the vulnerability of LLMs to structured interventions grounded in the polysemantic representations of two small models using SAEs. We show that model behavior can be steered toward specific feature directions by manipulating semantically unrelated yet interfering features via three intervention methods. Interventions distilled from polysemantic structures shared across the small models transfer to larger, black-box instruction-tuned models, indicating a stable and transferable polysemantic topology that persists across architectures and training regimes. Post-hoc annotation suggests that fewer than $30\%$ of these shared interference structures align with higher-order relations readily intelligible to humans; many counterintuitive cases may therefore serve as generators of testable hypotheses about latent knowledge structure. Finally, by leveraging the uneven distribution of features across neurons, we assess models’ sensitivity to neuron-level manipulations across degrees of polysemanticity and reveal asymmetric effects in “super-neurons.” Together, these findings provide a foundation for future work on the structural properties, vulnerabilities, and representational robustness of LLMs.
@@ Appendix A Related Work
A.1 A Brief Review on LLM Adversarial Interventions
@@ A.2 A Brief Review on SAE-based Intervention Techniques in LLMs
SAE-based interventions represent a promising direction for developing more interpretable and controllable LLMs. Recent research have introduced a diverse set of SAE-based techniques, such as clamping, patching, and causal tracing, applied across a range of use cases (Farrell et al., 2024; Cunningham et al., 2023; Marks et al., 2024). Empirical results indicate that these methods can be highly effective. For example, targeted unlearning via SAE features has been shown to suppress undesired capabilities with fewer side effects than global fine-tuning (Khoriaty et al., 2025; Muhamed et al., 2025), while feature-level steering enables more nuanced output control than prompt-based methods alone (Rajamanoharan et al., 2024). A key advantage of SAE-based approaches is their efficiency at inference time: they often require only a forward pass with lightweight vector operations and typically do not require model retraining, making them well-suited for real-time interventions.
@@ Appendix B Impact Statement
This work systematically investigates a semantic vulnerability in LLMs rooted in polysemanticity—where single neurons encode multiple semantically dissimilar features. We introduce four complementary approaches that expose this vulnerability: manipulating SAE-derived features, token gradients, and prompts to steer model outputs via semantically unrelated inputs, and intervening at the neuron level to reveal a correlation between polysemanticity and output sensitivity. We also identify a class of “super-neurons” whose amplification disproportionately alters model behavior, while masking them has a limited effect. These findings not only highlight the unique characteristics of the structural fragility of LLMs but also provide practical tools for probing and controlling their internal mechanisms. Our work lays a foundation for future research in AI safety and mechanistic interpretability, not only enabling defenses against such vulnerabilities and more targeted interventions for alignment, but also offering a theoretical lens into the model’s internal organization, revealing stable yet counterintuitive interference patterns that may reflect a form of unconscious knowledge association.
Appendix C The Use of Large Language Models (LLMs)
LLMs play a significant role in our study. We list the usages of LLMs below.
Synthetic data generation.
For the dataset used for intervention experiments, we use DeepSeek-V3 to generate incomplete sentences for next-token completion. Details are elaborated in Appendix E.
SAE feature auto-interpretation.
The SAE feature glosses in the main experiments are generated by GPT-4-mini. For cross-validation, we also annotate a subsample of SAE features using DeepSeek-V3. Details in Appendix J.
Feature pair association analysis.
To investigate whether the seemingly unrelated interference feature pairs present higher-order associations that are still comprehensible, we ask GPT-5-mini and DeepSeek-V3 to label these feature pairs for a post-hoc check. Details in Appendix M.
Grammar check for paper writing.
We use LLMs to refine phrasing, correct grammar, and improve readability during manuscript preparation, while all substantive ideas, analyses, and conclusions remain our own.
Appendix D Sparse Autoencoder Training
SAEs are a rapidly developing tool for probing the polysemantic structure of neurons (Shu et al., 2025). Given the activation vector $\mathbf{a}\in\mathbb{R}^{d_{\text{embed}}}$ from a particular model layer, an SAE projects it into a higher-dimensional sparse code $\mathbf{f}\in\mathbb{R}^{d_{\text{sae}}}$ in order to disentangle the multiple semantics that a single neuron may simultaneously encode. The forward computation and the resulting feature definition $\mathbf{f}$ are shown below: